[title Sol Summary – December 30th]
Summary Title: Martian Explorers
Author’s name: Pat Pesa
Mission Status: Crew is feeling accomplished!
Sol Activity Summary: Crew 218 had many sources of excitement today. We had two great EVAs that allowed for significant progress in progress in the Climate Monitoring, Detecting Radio Signals, EVA suit ergonomics, as well as starting the Geologist’s project of analyzing Sub-Surface Ground Structure.
Outside of our EVAs we also "solved" our water troubles (at least delayed them a day or two) by manually filling our Loft tank via human chain and buckets of water from the static tank (fireman’s style).
Look Ahead Plan: More EVAs! Also tomorrow will be New Years Eve, so we have some fun activities planned for the evening.
Anomalies in work: None
Weather: Partly Cloudy, around 32 degrees
Crew physical status: Healthy
EVA: EVA #4 this morning went to Galileo Road to take Climate data. EVA #5 this afternoon went nearly to the crossing of Brahe Highway (we had to stop slightly before due to Rover Battery levels), taking continuous Radio Signal Measurements, and a Seismic Survey for the Ground Structure Project.
Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report, EVA request, science report
Support Requested: none